OpenStudy (anonymous):

Assume f(x)=2x+8 and g(x)=5. What is the value of (f+g)(3). Assume f(x)= -2x+8 and g(x)=5x. What is the value of (f o g)(3). It'd be great if I could get some help!

OpenStudy (solomonzelman):

f(x)=2x+8 f(3)= 2(3)+8=6+8=11 --> f(3)=11 g(x)=5 ---> g(3)=5 (f+g)(3)= f(3) + g(3) = 11 + 5 =16

OpenStudy (anonymous):

@SolomonZelman thank you. Is that just for the first one? Would I follow the same process for the second or?

OpenStudy (solomonzelman):

No, fog is not adding them, I forgot what that is, I think you are to multiply what they are equal to.

OpenStudy (the_fizicx99):

The first one you're adding them, the second is -2x and you're multiplying

OpenStudy (the_fizicx99):

The second also has a 5x so the g(3) will be 15

OpenStudy (anonymous):

For the second, my answer choices are 17 3x+8 -22 29 It makes sense that it'd be 15 based off what you said, but it's not one of the choices

OpenStudy (the_fizicx99):

(f o g)(3) = -2(5(3)) + 8 f o g = -2(15) + 8 --> -30 + 8 = -22 (f o g)(3) = -22

OpenStudy (anonymous):

You are a lifesaver! Thank you for writing it out too. Could you help me with a similar problem? @tHe_FiZiCx99 Let h(x)=(g o f)(x)=1/(x+3)^2 What is the decomposition of h(x)?

OpenStudy (the_fizicx99):

f(x)=x+3 and g(x)=1/x^2

OpenStudy (anonymous):

thanks :)

OpenStudy (the_fizicx99):

:3 np

OpenStudy (anonymous):

Would it be bad if I asked for help on one more?

OpenStudy (the_fizicx99):

One more? Sure i gtg soon

OpenStudy (anonymous):

Okay! f(x)=x^2+3x+5 what is f(a+h)

OpenStudy (the_fizicx99):

That's only f(x)

OpenStudy (anonymous):

Hmm? The choices are these:

OpenStudy (the_fizicx99):

Can you post it with the question as well ^^

OpenStudy (anonymous):

Yeah. Here it is

OpenStudy (the_fizicx99):

Omg this is such a weird question, it would be perfect if A didn't have 5(a+h) because 5 is just a constant, I don't get why B has a 2ah, C seem interesting buts its being multiplied by x^2 + 3x and 5 I'm leaning towards D because its 3a + 3h + 5 but the x is only h^2, I think it should've been (a+h)^2

OpenStudy (anonymous):

Yeah it is a weird question. I tried to figure it out, but psh no luck there. Have a great night!